OpenStudy (anonymous):

Determine the solution to f(x) = g(x) using the following system of equations: f(x) = 3x + 12 g(x) = −9.5x − 13 x = −2 x = −3 x = 4 x = 5

OpenStudy (mathmale):

If f(x) = g(x), then you must equate 3x+12 and -9.5x-13. Do that now, please. Next, solve the resulting equation.

OpenStudy (mathmale):

Where's your new equation?

OpenStudy (anonymous):

wait do what?

OpenStudy (mathmale):

you must equate 3x+12 and -9.5x-13. Do that now, please. Next, solve the resulting equation.

OpenStudy (unklerhaukus):

f(x) = 3x + 12 g(x) = -9.5x - 13 & f(x) = g(x) SO 3x + 12 = -9.5x - 13 Solve for x.

OpenStudy (shadowreaper810):

f(x)=3x+12 g(x)=-9.5x-13 given that f(x)=g(x) so substitute both of them with the given values.. Now, 3x+12=-9.5x-13. solve for x